The Planet Altair IV is 16 light years from Earth. Twins Anna and Otto are 20 years old when Otto, a quantum mechanic on the United Planets Cruiser C-57D, leaves for Altair IV. Otto travels at 0.82c, quickly turns around and returns to Earth at 0.82c. a) How old will Otto be when he gets back? b) How old will Anna be when Otto gets back?
Answer #1

Total distance traveled in going and coming back is (16+16) = 32 light years at a speed 0.82C .

So, dt= 32 light years/ 0.82C = 39.02years

Again dt = dt0/√(1 - v2/C2​​​​​​)

So, dt0 = dt √(1- V2/C2 )

So, dt0 = 39.02 √ [1- (0.82)2 ] = 22.33 years

(a) Old of otto when he will back = (20+dt0) = (20+ 22.33) = 42.33 years

(b) old of Anna when otto will back = (20+ dt) = (20+39.02) = 59.02 years